OpenStudy (anonymous):

Two numbers, 3 and a, have a geometric mean of 9. Find the value of a.

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

Geometric mean of two positive real numbers a and b is defined by square root of (a*b) Hence we have square root(3*a)=9 this implies 3a=9^2=81 Hence a=81/3=27

OpenStudy (anonymous):

@omg1help

OpenStudy (anonymous):

ok so a=27

OpenStudy (anonymous):

yeah